OpenStudy (anonymous):

Determine the factors of x2 − 12x − 20. (x − 2)(x + 10) (x − 10)(x + 2) (x − 5)(x + 4) Prime

zepdrix (zepdrix):

We want `factors` of -20 that will `add` to give us -12. \[\large -5 \cdot 4=-20\]But when we try to add these,\[\large -5+4=-1 \qquad \qquad -1\ne-12\]Hmm so those factors don't work. \[\large -10\cdot2=-20\]Let's try adding them.\[\large -10+2=-8\qquad\qquad -8\ne-12\] You could also try \(-2\cdot10\) but you would still run into the same problem. This appears to be prime. Hopefully this helps you understand why.

OpenStudy (anonymous):

Thank youuu! It did help (: